OpenStudy (anonymous):

what is the equation of a circle wit diameter AB that has endpoint A(0,0) and B(8,6) a.(x-3)^2+(y-4)=5 b. (x-4)^2+(Y-3)^2=25 c.(x-4)^2+(y-3)^2=5 d.(x-3)^2+(y-4)^2=25

OpenStudy (radar):

That diameter is a line whose equation is y=3/4 x + 0 The midpoint is P(4,3) of the line and also the center of the circle. The radius would be:\[\sqrt{3^{2}+4^{2}}=5\] Do you have enough data now to complete the equation for the circle?

OpenStudy (radar):

For a circle whose center is located at P(4,3) then h=4, and k=3 and radius 5 would have an equation:\[(x-4)^{2}+(y-3)^{2}=25\] the 25 is the radius squared. working that equation would give you:\[x ^{2}-8x+16+y ^{2}-6y+9=25\] Expanding and simplifying we get \[x^{2}+y ^{2}-8x-6y=0\]

OpenStudy (radar):

I think!